OpenStudy (anonymous):

Ms. Soar flew 220 mi in 1.6 h. She then landed at the Aspen Airport to visit a friend. In the second portion of her trip, Ms. Soar flew 312 mi, which took 2.2 h. What was her plane’s average speed for the entire trip?

OpenStudy (anonymous):

I would like to know how this is solved...not exactly the answer..

OpenStudy (anonymous):

@K12awesomeness

OpenStudy (k12awesomeness):

What grade are u in?

OpenStudy (anonymous):

6th btw this is a k12 question

OpenStudy (k12awesomeness):

I'm in 6th also, so this should be easy. And I'm in K12 too.

OpenStudy (anonymous):

lol

OpenStudy (k12awesomeness):

Give me a sec, and I'll help you. :)

OpenStudy (anonymous):

k

OpenStudy (k12awesomeness):

137.5 + 141.8 = 279.3/2 = 139.65 Answer 139.65 miles per hour ;) So, you have to round... And that would be 140 miles per hour in all

OpenStudy (k12awesomeness):

I hope this helps you.

OpenStudy (anonymous):

wow...cool!!!! your good!!!:)

OpenStudy (anonymous):

i gave you a medal

OpenStudy (k12awesomeness):

thx, just tag me if you need me...

OpenStudy (anonymous):

how do i do that

OpenStudy (k12awesomeness):

like you did up above... @K12awesomeness

OpenStudy (anonymous):

ohhhhhhhh ya thanks

OpenStudy (k12awesomeness):

lol.. your welcome!!